Mr.Conley's present age is three times his daughter's present age. Five years ago, Mr.Conley was four times as old as his daughter was then. How old are they now?
Answer by ewatrrr(24785)


Hi
Mr.Conley's present age is three times his daughter's present age
Let x and 3x represent the daughter's and father's ages NOW respectively
Question states***
Five years ago, Mr.Conley was four times as old as his daughter was then
3x - 5 = 4(x-5)
Solving for x
3x - 5 = 4x - 20
15 = x, the daughter's age NOW. Mr Conley is 45
CHECKING our Answer***
40 = 4*10